Part A Starting at t 0s, a horizontal net force F (0.280 N/s )t i + (-0460 N/s2 )t2 j is applied to a box that has an initial momentum What is the momentum of the box at t 2.00 s? Enter the and y components of the momentum separated by a comma. P (-2.95 kg m/s)i(4.05 kg.m/s )j A2op p,, py kg m/s
Answer #1

Vector equation is given by

\frac{\mathrm{d} \vec{p} }{\mathrm{d} t}=\vec{F}

\vec{p}-\vec{p} _i=\int_0^t \vec{F} ~dt

\Rightarrow \vec{p}=\vec{p} _i+\int_0^t \vec{F} ~dt

\Rightarrow \vec{p}=\vec{p} _i+\int_0^t (0.28t\hat{i}-0.46t^2 \hat{j}) ~dt

\Rightarrow \vec{p}(t) =-2. 95\hat{i}+4.05\hat{j}+0.14t^2 \hat{i}-0.153t^3 \hat{j}

So, at t = 2 sec,

p_x=-2. 95+(0.14\times 2^2) =-2. 39~kg.m/s

p_y=4.05-(0.153\times 2^3) =2.826~kg.m/s